Instruments & Operations



Basic Guide to Operating the LBC

LBC User Manual:

Pages External to the User Manual:

Starting up LBC

The LBC Control system operates within a firefox browser. To bring it up,

Back to Top

Executing an OB

  1. Go to the "OB Execution" Page: To execute an OB, go to the "OB Execution" page. As the image shows, this page is divided into 8 sections. Along the top row, the "OB Operations" control is at the left and "OB Info" is displayed in the upper right; From left to right along the middle row: "OB Progress" is shown at the left, the guiding offsets set to the telescope are shown on the right (rotator guiding is not enabled but RA and DEC offsets will be shown) and the middle section displays weather information Finally along the bottom: the left panel displays telescope information and the right two panels display the status of the blue and red channels.

  2. Load the OB: To load an OB, click on the "Browse..." button in the "OB Operations" section. Select the OB you wish to run and click the downward pointing arrow to upload it. Or, to create an OB, click on the "Create an OB" button within the same section. A new window will pop up and after you have finished filling in the fields in it, click the downward pointing arrow at the bottom of it to upload it to the OB execution page. This newly created OB will be called "fast". The "Create an OB" functionality is discussed in detail in the OB Preparation page. When the OB is uploaded, the downward pointing arrow in the "OB Operations" section should change from black to grey and the "play", "pause" and "stop" buttons (right-pointing, double vertical dashes, box) should turn from grey to black. The OB name, target name, type and coordinates will be displayed in the "OB Info" section. If the observation should be carried out at the current telescope pointing, then current will be written instead of the coordinates. See the OB Preparation page for information on how to indicate that the current pointing should be used.
  3. Review Settings: Before executing the OB, review the settings for Telescope, Guiding, Channel, Mirror and Scaling.
    • A check mark to the left of Telescope indicates that:
      • the telescope will preset (slew) to the coordinates in the OB (or remain where it is if the Coords are listed as current),
      • even if "Preset LBT" is indicated in OB, the telescope will not slew unless Telescope is checked in user interface
      • track the target at the sidereal rate according to the trajectory calculated by PCS and
      • the LBC rotator will track (rotate) according to the trajectory calculated by PCS.
    • if "Rotator Tracking" is not checked in the OB, but Telescope is checked on the User Interface, the rotator will track
    • If Guiding is checked, then guiding offsets (corrections to tracking) measured from the technical chip images will be sent to the telescope and used. Leaving it unchecked will not prevent technical chip images from being taken and offsets computed, only no corrections will be sent to the telescope.
    • A check mark to the left of Channel indicates whether to use the Blue, Red or both channels; and
    • A check mark to the left of Mirror indicates that the instrument will be able to command adjustments to the primary mirror position (necessary for focusing).
    • Insure that the exposure time scaling factor is what you want (defaults to 1.00).
  4. Executing the OB: To execute the OB, click the play button. First the telescope slews to the desired coordinates, then the mirror is adjusted for focus and collimation. During this time (seconds to a minute or two), the status written below whichever channel was checked will remain "ready". Once the mirror is in position and the filter and rotator are being moved, the status is shown as "preset/read/save". Look at the Log Analyzer page if you wish to find out just what is happening at the time. The OB will run to completion unless the stop or pause buttons are pressed.
  5. Stopping an OB: Pressing the stop button will interrupt the current exposure or readout and will NOT save it. Pressing play will resume the OB, and retake this exposure. Importing the OB again before pressing play will start the OB from the beginning.
  6. Pausing an OB: Pressing the pause button will cause the OB to pause after the current exposure finishes. This exposure will be written to disk. When you resume the OB, the next exposure in the sequence will be taken. The exposure time scaling can be adjusted when the OB is paused; this may be helpful when taking flat fields.
A note on Exposure Times and Total OB Duration: Back to Top

Looking at the Data

Images from both the science and tracker chips are sent first to the read-only /newdata directory, which is mounted on the observer workstations. From there the data are copied to a UTdate (YYYYMMDD) subdirectory of /Repository, which is also mounted on these workstations and read-only.

Both /newdata and /Repository are NFS-mounted and read-only. Therefore, quick analysis is most efficiently done in a subdirectory of the local directory, /scratch.

Examining your data real-time

There are several ways to examine your images as they arrive in the newdata:


LBTtools is an IRAF package which contains tasks useful to quick analysis of data taken at LBT. It should be available from all partner accounts.

Setting up LBTtools:

LBTtools tasks often used with LBC:

Back to Top

Beginning-of-Night Tasks

The telescope should be handed over to the night crew 2 hours before sunset. The operator and instrument support personnel will fill the LBC cryostats and will conduct a safety inspection to insure the telescope and instruments are ready. Once these are completed, the observers are ready to obtain calibrations: typically a set of biases well before twilight and then twilight flat fields. The OBs to obtain these are included in the Calib_OBs.tar package; although each partner should have a set of these already in their account.

Back to Top

Active Optics: Focus and Collimation

At the start of the night, after each slew, and about every 30 minutes while tracking a target, the telescope focus and collimation must be adjusted. This is done by taking and analyzing extra-focal pupil images. At the start of the night, the telescope operator should have "Cleared Active Optics" to remove any residual active optics corrections (this clears both mirror shape and collimation corrections) and added +1000nm of Z11 to SX (Blue). The procedure for focussing and collimating the mirrors follows. Since February 2009, the focus/collimation procedure has been automated through the IDL program "dofpia" which both takes and analyzes the extra-focal pupil images, iterating until convergence has been achieved.

To focus and collimate on a field

Back to Top

Shutting Down LBC

To shut down LBC at the end of the night, follow the steps below. Note the order, and wait until one step has finished before starting the next:

Back to Top

End-of-Night Tasks

Back to Top

Getting your Data

Data Flow

The images first are copied to the windows PCs (located in the upper left treehouse). From there they are copied to the /newdata directory, which is mounted on the observer workstations. On average it takes 6 seconds for images to arrive in /newdata. From /newdata they are copied to /Repository/UTDATE (where UTDATE is in the format YYYYMMDD) and to the archive in Tucson, which is mirrored in Italy and Germany.

On the mountain, images older than one day are removed from /newdata, so observers should access /newdata for current data and /Repository/YYYYMMDD for data from previous nights of their run.

Tech chip images and OB files are stored in /newdata for the current night and, for previous nights, in the subdirectories tec and ob of /Repository/YYYYMMDD.

Getting your image files

At the end of the night or run, the images can be copied directly from /Repository/YYYYMMDD to a storage medium: portable hard disk or DVDs. They can also be obtained through the LBT archive by pointing your browser to the URL,, and logging in using the information which you should have received from your LBT partner coordinator.

Back to Top


When a problem arises, check the "Log Analyzer" for error or warning messages. New notes and status information appear in the Log on timescales of seconds so you may need to select (via the check boxes) only the systems that are related to the error message and deselect the other systems. Click "parse" to refresh the view.

Camera or Trackers

Camera or trackers do not come up: There are several failure modes. These may occur during observing also. Check the log for warning or error messages.

LBC spontaneous Shutdowns: Suddenly the status of LBC reverts from "ready" to "enabled". Issuetrak #298 reports one such instance when there was a CCD Controller error. The only thing to do is restart LBC. Do not leave LBC in this state as the Windows machines and CCD Controllers are in an unknown state. Even if you are at the end of the night, LBC should be turned on and back off again. A close relative of this failure is the "hinibit" (inhibit) failure where the LBC safety software shuts down the camera when it fails to read the dewar temperature or vacuum 10 times in a row.

Blue Controller shuts off --- temperature safety limit exceeded: In late-Jan/early-Feb of 2008, there were problems with the Blue camera detector temperature sensor. When this gets a few bad readings, it makes an error that stops the Blue science camera. A power cycle is required to revive it (i.e. TurnOff/TurnOn). See Issue #1128 for details. The LBC software was fixed and the Blue Camera is reporting the correct detector temperature all of the time that the camera is on.

Shutter gets stuck: If the shutter sticks closed, the images will look like biases --- no stars visible and counts close to the bias levels. If the shutter sticks open, or partially open, then trails will be seen, in the +y direction, from all of the stars caused by the readout clocking. Fainter trails will be seen in the -y direction from the preflush clocking. (Depending on image scaling, you may only see these trails for the brightest stars). The image below was taken with LBC-Red during morning twilight, when the sky background was high. The image on the left shows a strong right to left gradient on chips 4 and 2 (chip 3 was out of service at the time). On the subsequent images, the other blade closed part way, leaving only a strip of aobut 1000 rows exposed to light, as the image on the right shows. The trail of a bright star about one-thirds of the way from the bottom of chip 2 extends in the positive Y direction and makes a kink as the telescope dithers during the sky flat sequence. The image gallery also contains an example image taken with when the shutter stuck was stuck all the way open.

Reinitializing the shutter by power cycling may restore functionality. This may be accomplished, by the observer, by stopping and restarting the LBC, i.e. on the User Interface, turn off and back on "Other Systems", or by the support astronomer through the testpower3 program. Give this a couple of tries, but if it does not help, then the support astronomer may try the low level commands in issuetrak #2532.

If none of these suggestions works, then the Instrument Specialist will need to be contacted, and an inspection of the shutter should be done. The LBC-Blue hub has a shutter toggle switch. This should be positioned to "Auto" for remote operations (control by the User Interface), but can be toggled to move the shutter manually. Try this and listen as the shutter moves. This type of switch does not exist (this needs to be confirmed) on LBC-Red.

Profibus Errors: Profibus is the serial bus protocol used to communicate between the CMU and the various encoders on the rotator and filter wheels in the camera.

Filter wheel errors: Occasionally a filter move stops with an error. There are two common problems: check the LogAnalyzer to see the error message.

Images appear trailed:

Related issuetraks are: #2045 and others.

LBC User Interface Hangs: This happens sometimes and comes in several flavors. Sometimes the UI browser seems to be hung waiting for the CMU, while other times the UI seems to be refreshing without updating the important parameters related to the present observation. Sometimes things are happening and just not being reported: if the mirrors have stopped moving and the telescope has arrived at position, both of these are reported by the IIFGUI, then start looking for the cause of the hang elsewhere. This problem is documented in Issuetrak #180. More recently, we have seen the UI hang while shutting down after another error has occurred. The system which had the error hangs in the state uninitializing. In this case, it is necessary to run and (Number 5 from the list below). The following actions, which are listed from least to most disruptive of observing, are recommended to restore full functionality:

  1. Try changing from the OB Execution page to the Power Control page and back. This will not interupt the on-going observation.
  2. Try simply waiting until the end of the on-going exposure. Often the UI will recover when the image begins to readout.
  3. Try killing your Firefox browser window on the local workstation and starting a new one. This will not interrupt the on-going observation.
  4. If LBC loses contact with TCS during a preset, it will wait 5 minutes before timing out. Always wait at least 5 minutes before proceeding to more drastic action. You should also spend the 5 minutes trying to understand why the TCS failed to respond (e.g. there could also be a TCS shared memory problem or other telescope issue).
  5. Have a root user log into CMU and run the and scripts in /home/lbccontrol/ (see below). This requires going through the full power-up cycle (10 minutes) on the LBC power control webpage. Any observation in progress will be lost.
  6. Have a root user log into CMU and reboot (see detailed instructions below). This also requires going through the full power-up cycle on the LBC power control webpage.

Main Mirror Fails to Focus: When the temperature gets down to -5 degC, the SX Primary Mirror gets a bit sticky in its hardpoint motions and sometimes fails to meet the position tolerance. When you get the LBC pop-up window that says "Left Main Mirror Fails to Focus", the observer should click "OK" in the LBC pop-up window and allow the OB to continue. Then it should be verfied on the PSF GUI that the mirror is only out of tolerance by 10 microns or so. Compare the "total collimation" X Y Z with the "platform position" X Y Z. (remember that X Y Z are in millimeters and RX RY RZ are in arcsec) This check is to make sure that you aren't observing with some actual mirror support failure which would give similar symptoms except with position errors of hundreds of microns or more. This problem is documented in Issuetrak #383.

Missing images: There are several reasons an image may be missing from the Repository, lbcarchive or lbtarchive:

General instructions and information:

Rebooting the CMU: Before rebooting the CMU, turn off the camera (science and tracker cameras and shutters, rotators, filter wheels, housekeeping) as far as possible. If the camera or interface is in a 'hung' state, then there is nothing that can be done. Though not necessary, it is a good practice to close the user interface also. Then, to reboot the CMU, open an xterm and “ssh" as root into the CMU (lbccontrol): ssh root@lbccontrol , type 'who' to check that there are no remote users and if not, or once they are suitably warned, type "reboot". At this point, the xterm will hang, and once the machine has rebooted, it will return the prompt of the workstation to which you are logged in (e.g. obs3,5 or 4). The reboot will take about 3 minutes. Once the CMU is back up, open the user interface and turn back on the “housekeeping” and “other systems”, and finally “connect LBT”.

Connection to CMU lost?: Sometimes the connection to the CMU is lost. The machine may have crashed or been powered off. Ask the operator or instrument scientist (call the support astronomer if you need assistance) to go into the treehouse to press the reset button on the CMU.

The commands, and

Back to Top

Engineering Information (for support staff)

Filter Positions (Dec-2012)

Blue Filter Wheel Positions (2011 January)
Wheel #1Wheel #2
positionstepsname positionstepsname
19676 115200
232204U-BESSEL 237728r-SLOAN
354732V-BESSEL 360256empty
477260SDT_Uspec 482784(was pinhole)
599788B-BESSEL 5105312g-SLOAN
Red Filter Wheel Positions (2012 December)
Wheel #1Wheel #2
positionstepsname positionstepsname
171931 18825
298282V-BESSEL 235469r-SLOAN
3124633I-BESSEL 361952i-SLOAN
419229R-BESSEL 488271F972N20
545580Y-FAN 5114848z-SLOAN

Back to top of the Engineering Information Section

Making new Collimation Models for the LBCs

Periodically, the Collimation Models for the LBC need to be updated. Ideally this should be done once a month - during the technical nights which precede an LBC run and after changes to the LBC hub (moving L2 or moving the entire hub as is done twice a year because of seasonal temperature variations in the telescope and swing arm structure).

Data for the collimation model should be obtained in conditions of good seeing (less than ~1"). The procedure for making a collimation model is basically (1) to collimate (run dofpia on star fields at a number of different elevations), (2) note the final pupil image at which dofpia converged, (3) make a list of the mirror position: X,Y,Z,RX,RY; as well as elevation and temperature for each of these final images, (4) run an IDL script to make a quadratic fit to these data and output the coefficients to a file, and (5) copy this file to the telescope.

A step-by-step guide to generating a collimation lookup table follows:

  1. Ask the operator to "Clear Active Optics" and add Z11 = +1000nm to SX.
  2. Slew to a star field and run dofpia to focus and collimate. If this is the first focus/collimation of the night, it is best to slew to a moderate elevation. Something along the meridian and with elevation ~70deg would work well. Coordinates need to be given in RA and DEC, but the slew can be done either by the operator or by the observer, through the 'fast OB' interface.
  3. A cautionary note about mirror travel limits: Before starting to collect data for the collimation model, insure that the mirror travel limits (+/- 2.6 mm) are not hit at very high (greater than 80 deg) or very low (30-40 deg) elevations. The PSF GUI will display the mirror position in red when a limit is hit. The limits tend to be hit at high/low elevations on cold nights (sub zero?) during which it is not possible to focus and collimate over the entire elevation range using the same set of global or pointing offsets. To move away from an X or Y limit, you must make a coma-free pointing correction . Moving away from a Z limit requires the entire LBC hub to be shifted, which is normally not a nighttime operation. The telescope operator can move away from an X or Y limit via the PSF Arbitrator Control GUI which is brought up by clicking the Arbitrator button on the Left PSF Control GUI. This will change the values for X/RY or Y/RX that are shown under Pointing Offsets in the PSF Primary Mirror Active Optics and Collimation Control. Note these changes since they are needed for the collimation model. The data interpretation will be much easier if you obtain all of the collimation data with the same values for the global offsets and pointing corrections. Therefore, it is behooves you to choose a good compromise set at the beginning.
  4. Run dofpia at this position.
  5. Repeat steps 1, 2 and 4 to sample well the range of elevation. Usually I aim to step through the 30-90 degree elevation range in 10 degree increments and to cover the entire range twice, once going up in elevation and once going down.
  6. Make a list the final pupil images which dofpia reported as collimated and after which it exited.
  7. Run the task pcollim which will take this list and output a new list of header values for the mirror position (MIRRORX,MIRRORY,MIRRORZ,MIRRORRX,MIRRORRY), telescope elevation (TELALT) and ambient temperature (LBTTEMP). The task must be loaded by task pcollim = /home/okuhn/reduce/
  8. Run the IDL program to fit the table of data values listed in the output of pcollim.
    • cd /home/lbcobs/Collimation/new/
    • idl
    • mkcoll_lbc, "input filename" (where the name of the input file should be in quotes)
    • cp Collimation_lbc.dat {SX,DX}PMLBCCollimation_.dat (The IDL program's output is always named "Collimation.dat" so it must be copied to avoid overwriting it on the next run). The format with UT date is strongly suggested.
  9. Finally, copy new Collimation table to the telescope account on obs1.
    scp {SX,DX}PMLBCCollimation..dat telescope@obs1:/home/telescope/TCS/Configuration/PSF/new/
    and then copy it to the working name (i.e. without the date), {SX,DX}PMLBCCollimation.dat
  10. Ask the operator to click the button "Initialize" on the PSF GUI; this will load the new collimation table.

Back to top of the Engineering Information Section

Focus Offsets

On nights of stable sub-arcsecond seeing, we have measured the relative filter focus offsets. First we focus and collimate, then we run an OB which steps the primary mirror by 0.04 mm at 5 positions centered around the best-focus. Usually only chip 2 is needed, but obtaining data with all chips allows a study of focal plane curvature which is useful in optimizing the position of L2.

How to measure filter focus offsets: Ideally these measurements should be done in good seeing (0.7 arcsec) and thermally stable conditions. The procedure should be the following:

  1. Use the IDL procedure dofpia to focus and collimate.
  2. Use superfoc OBs to take FOCUS sequences of 7 images separated in focus by 0.060 mm for each filter. These OBs can be found in /home/LBTO/Calib_OBs/SUPERFOC/ (the Calib_OBs.tar package is available at the link).
  3. Analyze the superfoc sequences:
    • run the LBTtools task foclist after the sequence has finished. foclist makes a list of the last n focus images in Repository (or whatever data directory is indicated in LBTtools) that were taken through the specified filter, and appends the designation for the extension which refers to the middle chip, LBCCHIP2. (Note: if foclist is not run immediately after the focus sequences of interest, but after more have been taken, it will select the later images instead of those from the sequence of interest.)
    • Check the output of foclist to confirm that the FOCSEQN number is the same for all of the images
    • use the IRAF task "starfocus" (the command syntax is displayed as output of foclist) to determine the best focus.
      • select stars by positioning the cursor over them and typing "m"
      • to quit from the star selection and view the fit result, type "q"
      • Note: try to select stars always from the same area of the chip. I usually zoom in to the center or the rotator center of CHIP2 and select about 5-7 stars.
  4. Iterate steps (1) - (3), re-focussing and re-collimating with 'dofpia' sequences every 20 minutes. Continue until you get consistent results.
  5. When confident of the focus offset, update the file leftchannelfocus.dat or rightchannelfocus.dat in /home/lbccontrol/conf/ with the new filter focus offsets. Use "vi" or "mc" (Midnight Commander) to edit the files.
  6. Take short "object" images in each filter to check that the offsets are applied correctly.
Changes in the filter offsets need to be updated in:

Focus offsets between the LBC-Blue Filters
Date offset[mm] made after dofpia completes
28-Jan-07 0 0.05 0.32 - 0.07 0.08
17-Mar-071 0 0.072 0.32 0.4862 0.033 0.044
10-Jun-073 0 0.053 0.353 0.454 0.086 0.083
02-Jun-09 0.435
15-Apr-104 0.07 0.068
05-May-105 0 0.046 0.344 0.446 0.091 0.065
29-Dec-106 0.07 0.116 0.414 0.516 0.161 0.135
9-Mar-117 -0.02 0.026 0.324 0.426 0.071 0.045
19-Sep-118 -0.02 0.026 0.3 0.414 0.04 0.035
30-Sep-159 0.00 0.045 0.31 0.43 0.068 0.06
1.Better measurements, but their V-zeropoint may be less consistent. We believe that the B filter was inverted during cleaning on 9-Mar-07.
2. updated to 0.450 on 11-May-07
3. These had the V-zeropoint adjusted on 9-Jun-07. (JMH)
4. First new Blue Focus offsets after the Z22 Zero Point was corrected
5. On the 05-May-2010 technical night, all of the Blue Filter focus offsets were redetermined by M. Pedani and J. Hill. These are with the corrected Z22 Zero Point from April.
6. An +0.07mm offset was added to all filters, based on 3 measurements of the V-BESSEL best-focus after the BlueZ22Zero was changed from 1.6 to 2.3.
7. A -0.09 offset was added to all filters, after the BlueZ22Zero was changed back to 1.6, SphOff changed from 88 to 338, BlueFinalZ11 changed from 400 to -100 and FocusAlphaBlue changed from 78 to 78.5. The TWiki Log for 9 March 2011 provides more information and rationale for these changes to the lbcfpia configuration file.
8. Current offsets as of 13-Oct-2014, which include updates made on 2011-09-19.
9. The use of a different region of chip2 by dofpia occasioned remeasuring focus offsets for at least the reference filters, and for LBCB, for all filters. Updates are based on data taken on 30-Sep-2015.

The focus offsets for the red filters were measured at the beginning of 2008. The most recent values in the configuration file are listed here.

Focus offsets between the LBC-Red Filters
Dateoffset[mm] made after dofpia completes
V-Bessel R-Bessel I-Bessel r-SLOAN i-SLOAN z-SLOAN Y-FAN F972N20 TiO_784 CN_817
before 30-Dec-07 0.13
31-Dec-07 0.121 0.181 0.141
13-Jan-08 0.302 0.2052 0.102 0.09
04-Mar-08 0.0963
09-Feb-08 0.024
09-Mar-08 0.22 0.045 0.026 0.08 0.04
10-Mar-08 0.0
28-May-08 -0.0082 0.0252
30-May-08 0.00
21-Jan-09 0.2393 0.1043 0.073 0.1613 0.0893 0.0763 0.0813 0.223
20-Mar-095 0.206 0.1297 0.0788 0.1399 0.0810 0.0611 0.06312 0.2213
21-Mar-095 0.114 0.02215
15&16-Jun-0916 0.16 0.03 -0.02 0.057 -0.005 -0.01 -0.037 0.18
17-Sep-0917 0.08
15-Apr-10 0.08 0.123
05-May-1018 0.079 0.013 -0.047 0.046 -0.021 -0.023 -0.026 0.139
26-Mar-1119 0.069 0.003 -0.037 0.036 -0.031 -0.033 -0.036 0.129
19-Sep-1120 0.069 0.015 -0.046 0.040 -0.01 -0.035 -0.018 0.129
30-Sep-1521 0.050 0.005 -0.04 0.025 -0.015 -0.03 -0.04 0.11 0.025 0.025
1. rough estimate
2. pretty good fit
3. good fit
4. Measured with reference to the intra-focal best-focus and the extra-intra offset of 0.214mm added.
5. Good seeing and stable. Seeing around 0.5". Nevertheless, there was much scatter in best-focus measurements.
6. Average of 3 measurements: 0.1912,0.1942,0.2189.
7. Two measures at 0.125 limit of OB focus range. Edited OB and got 0.1292, but later in the night got 0.110. 0.129 is consistent with the two previous measuresso used it only and neglected 0.110 measurement.
8. Average of 4 measurements: 0.0627, 0.0886, 0.0880, 0.0726.
9. Average of 2 measurements: 0.01398, 0.01382 (Collimation done in between but the measurements were not separated by large time.)
10. Average of 3 measurements not near limit: 0.1154(near limit), 0.0973, 0.0769, 0.0669.
11. Best focus varied a lot over the chip on both sequences: sequence one results are 0.0739 or 0.0689 and sequence two, 0.0405, 0.0571. Adopt 0.06.
12. Average of two measurements: 0.0590, 0.0679.
13. Average of two measurements: 0.2313, 0.2162.
14. This R-Bessel measurement of 0.1 is closer to the earlier value and last nights' end of night value (0.11, remeasurement gives 0.1),so I restored R-Bessel focus offset in LBC filter focus configuration file to 0.104. Comparing 20-March to 21-Jan results, R-Bessel was the only one that I increased; the rest dropped by a small amount, about 0.02mm. Restoring R-Bessel is more consistent with the general changes from January to March.
15. This z measurement just highlights the scatter we are seeing in z. I do not know the cause; leaving at 0.06
16. These were measured on the nights of 15 and 16 June 2009 after the Z22 zeropoints ({Blue/Red}Z22Zero in the lbcfpia.cfg file) were adjusted.
17. Stars through V-BESSEL with previous focus offset (0.16) were donuts. One measurement on 20090911 gave ~0.10 instead, and a second measurement done on 20090917 gave 0.08. Replaced 0.16 with 0.08 and got well-focussed images.
18. Focus offsets redetermined on 05-May-2010 by M. Pedani and J. Hill.
19. Subtracted 0.01mm from all focus offsets based on r-SLOAN results on 20110309 and one Y-FAN result on 20110326. Mistakenly added 0.01 to I-BESSEL, but images looked good and left as is. Seeing at times was sub-arcsecond butdidn't seem stable enough to warrant further measurements on 20110326.
20. Status as of 2014-Oct-13, which includes updates made on 20110919 for R-BESSEL, I-BESSEL, r-SLOAN, i-SLOAN, z-SLOAN and Y-FAN.
21. Use of a different region of chip 2 by dofpia occasioned remeasuring at least the focus for the reference filters. For these a -0.02mm offset was indicated.For R-BESSEL (4), I-BESSEL (5), i-SLOAN (3) and z-SLOAN (2) there were at least two measurements (actual number is in parenthesis after the filter name), from which focus offsets relative to the reference r-SLOAN could be determined. For the other filters, the -0.02mm offset was applied.

If the Z position of the mirror hits a limit at the extra-focal position, you will need to obtain intra-focal pupils. lbcfpia will recognize these as such and will compute corrections. The corrections that are output for focus (Z4), and coma (Z7 and Z8) can be used as they are, but the sign of the correction for spherical (Z11) is flipped. The corrections for astigmatism are not easily converted. Ask the telescope operator to input the output values for Z4,Z7,Z8 and -Z11. There is also a focus offset that needs to be entered when using an intra-focal pupil image to focus and collimate. For the Red channel, this was estimated to be -0.214mm. This offset can be accounted for either by subtracting 0.214mm from all of the filter focus offsets that were referenced to focus obtained from extra-focal pupils (there are files called rightchannelfocus_intra.dat and rightchannelfocus_extra.dat) or by asking the telescope operator to subtract 0.214 from the default global Z offset for DX.

Focus offsets between the LBC-Red Filters wrt intra-focal pupils
Filteroffset (mm)Datecomments
V-BESSEL 0.086 13-Jan-2008quite good
R-BESSEL-0.009 13-Jan-2008quite good
I-BESSEL-0.118 4-Mar-2008measured, quite good
r-SLOAN -0.0231-Dec-2007rough
i-SLOAN-0.07 31-Dec-2007rough
Y-FAN-0.1424-Mar-2008measured but not good fit

FilterPlatescale (arcsec/pixel)
1. Reported by Vincenzo Testa 14-May-2007

Back to top of the Engineering Information Section

Guiding parameters

Guiding parameters are set on the lbccontrol computer (root access only) in the files bluechannel.conf and redchannel.conf. For binocular operation with the Red trackers used to guide the telescope (Red as Master) and the Blue trackers used to determine corrections for the Blue channel (Blue as Slave), the parameters are as follows:

TypicalExposureTime 4 8
MaximumExposureTime 3264
MinimumScientificExposureTime 2432
LoopStartDelay 151151
MinGuidingCorrection 0.010.1
MaxTrackingErrorPerSec 0.200.20
MinimumTrackingStars 11
AzTrackingGain 0.90.9
ElTrackingGain 0.90.9
1. The LoopStartDelay, the time (sec) before the tracker image is taken, was increased from 10sec to 15sec in January to accomodate a temporary change to the controller software that was made to alleviate Blue Controller shutoff due to too many bad temparature readings.

Note that the image taken immediately after this file is edited and saved uses previous values; the changes are made only for the subsequent image and onwards. LBC does not need to be run down and back up to accept changes made to guide.dat. We shortened the TypicalExposureTime to 8sec in late December in order to better fight the problem of the elevation jumps.

The MinimumScientificExposureTime on the science array for which guiding with the technical chips will be activated is a parameter set in the bluechannel.dat file (root access only). This is currently (21-Dec-06) 24 seconds. LBC must be run down and back up to accept changes made to the bluechannel.dat file.

Back to top of the Engineering Information Section

Directions of Dither Offsets

Measured at position angle = 0:
Dither offsets in the +X direction move the star image right (+X) on Science chips 1-3, and down (-Y) on Tech chip 1.
Dither offsets in the +Y direction moves the image down (-Y) on Science chips 1-3, and left (-X) on Tech chip 1.

Tech chip 1 is right of science chip 1. Pixels are 13.5 microns, and the blue plate scale is 0.227 arcsec per pixel.
The dither offset onto Tech chip 1 is +860, 0 arcsec from the field center.
Back to top of the Engineering Information Section

Rotator center

1. The Blue cryostat was dismounted during the 2009 summer shutdown. But the pins should insure good alignment when remounting, and we have not measured such large (6-7 pix) differences since they were put in. 6-7 pix * 13.5mic/pix = ~90 microns. What is the specification for alignment; less than 90 microns? Headers will still use old (2008-09-22) values for the time being.
Back to top of the Engineering Information Section

Rotator velocity scale

A scale factor in rotator.dat will adjust the rotator velocity relative to the trajectory calculated by PCS. This should be 1.00.
Back to top of the Engineering Information Section

Reconfiguring LBC to store data locally rather than transfer images to the archive.

If the LBC archive is down and does not come back up after a reboot, as a last resort, you can take data and store it on the the CMU. You will need to login to the CMU and edit bluechannel.conf and lbc.conf. In bluechannel.conf, look for Camera.Upload.Address and Trackers.Upload.Address. These normally are set to to send data to the LBC archive. To send data instead to the CMU:/images/tftp change the "201" to "1". In lbc.conf, look for UploadAddress? . Again, change the IP address from to Remember to change these back once the LBC archive machine is up and running again!!!

Back to top of the Engineering Information Section

Pointing Model Data

Pre-requisites for collecting LBC pointing model data
  1. Have the operator reset the Az and El encoders before the pointing run.
  2. You need to have a good collimation lookup table for the primary mirror(s). You do not want to run active optics much in the middle of the pointing data. That could change the pointing, adding to the rms residuals of the model.
  3. You need to turn on the mount logging on PCS Pointing page. It logs continuously once on, every 2.5 seconds, so all you need to do is turn it off at the end of the night and copy the logfile to somewhere.
  4. You really only need to save data from chip2. That will also make it easier to download the data to Dave's laptop. The field of view is large enough that we can always do a decent WCS on the available stars. I need at least ~21 well-distributed fields to make a "low-order" pointing model (just mount and tube flexure). Use 10s exposures to make sure there are a couple of valid mount snapshots during the exposure. You can point to anywhere on the sky (it does not have to be on pointing stars), I'll derive the PM for a virtual star on the detector.

Back to top of the Engineering Information Section
Back to Top